# Detecting Dataset Bias in Medical AI: A Generalized and Modality-Agnostic Auditing Framework

### v1

@@ 4.2 Measuring Utility and Detectability
We use information theory and principles of causal inference to measure the utility and detectability of dataset attributes. Utility is measured as the mutual information, $MI(A;Y)=H(Y)-H(Y|A)$ after adjusting for chance as per [44, 45] and represents the reduction in uncertainty about the value of $Y$ by observing $A$ after adjusting for the entropy of the underlying distributions and chance based on the number of categories. We rely on the faithfulness assumption which implies that if there is a relationship between $A,Y$, then $MI(A;Y)>0$.
For detectability, the objective is to determine the ability to infer the value $A$ from the data and we accomplish this by first training a surrogate model $f:A\rightarrow X$ to predict $A$ from $X$. We next leverage clinician and domain expertise to establish the likely directions of dependency between data and labels. The primary role of this step is to understand whether the relationship is $Y\rightarrow X$ or $X\rightarrow Y$. Our ultimate goal is to identify whether there is a relationship $A\leftrightarrow X$ and if so measure its strength. However, we need to control for the potential of information leaking through $Y$ and this process relies on the assumed/known relationship between $Y,X$.
@@
In the causal case (i.e., $X\rightarrow Y$), we are able to directly estimate the relationship without conditioning. In fact, we cannot condition on $Y$ since $Y$ represents a collider (i.e., $Y$ may be dependent on both $X$ and $Y$). In that case, measuring the strength of relationship between $X\leftrightarrow A$ given $Y$ could give falsely inflated results since conditioning on a collider $Y$ creates an otherwise non-existent association between its parents ($X,A$). For example, if two conditions both increase mortality rate through different pathways, then given that a patient has died, knowledge of either attribute provides information about the other (e.g., given $A$ is the presence or absence of trauma wounds and $X$ contains information relating to disease, a patient who died but did not have trauma wounds is more likely to have had disease: $A\leftrightarrow X|Y$ even though $A\perp X$ ). As a result, we do not condition for cases where we expect $X\rightarrow Y$ and instead directly estimate $MI(A;\hat{A})$.
@@ 4.4 Bounding Performance Risk
While utility and detectability measures provide a quantitative measure of bias, how to interpret them in the context of potential model performance risks is less clear. The magnitudes of these measures are not easily comparable across datasets so we cannot rely on thresholds or conventions to directly assess risk. We could instead look to the relative rankings of attributes with respect to utility/detectability to understand attribute risk, but cannot directly translate values to drops in task-relevant metrics like AUC.
@@ 4.5.2 Stigmatizing Language in Electronic Health Records
We fine-tune BERT [10] models for both attribute prediction and each of the three clinical tasks: (1) compliance, (2) appearance and demeanor, and (3) credibility and obstinacy. Attributes we evaluate as potential shortcuts include patient race, gender, and clinical specialty (see Appendix A). The dataset consists of manually annotated samples with a context window of 10 words before and 10 after each identified potentially stigmatizing anchor word [19]. The full list of anchor words for each task is available in Appendix A.0.1. As each task has a unique set of anchors, the input and attribute metadata are different across tasks. Hyperparameters for all task and attribute models are held constant. We use AdamW [30] with a fixed learning rate of 5e-5 and weight decay of 1e-5, a batch size of 16, dropout with probability 0.1, 10 training epochs with early stopping, and class balanced cross-entropy loss for all experiments.
@@ 4.5.4 SPLIT method
As an alternate baseline form of estimating detectability and following Gichoya et al. [15], we test whether task model representations implicitly encode attribute information. Given a pre-trained classifier, we remove the final fully-connected layer and replace it with a randomly initialized linear layer. We freeze the weights of the pre-trained network and finetune only the linear layer to predict the specified attribute value. As before, we perform cross-validation and measure the performance of the finetuned model on the aggregated predictions across all folds. Better than chance performance is an indicator that model representations encode some degree of attribute information. However, it does not necessarily indicate that the dataset itself is biased as some attribute information may be relevant to solving the task even when the dataset is balanced with respect to the attribute itself.
@@ 5 Data Availability
Data for the ICU mortality prediction and skin lesion classification cases are publicly available. Because clinical notes from the Johns Hopkins Medicine (JHM) dataset contain identifiable information, they cannot be shared outside of our study team; however, all code to replicate the experiment on other datasets is available.

### v2

@@ 4.2 Measuring Utility and Detectability
We use information theory and principles of causal inference to measure the utility and detectability of dataset attributes. Utility is measured as the mutual information, $MI(A;Y)=H(Y)-H(Y|A)$. It represents the reduction in uncertainty about the value of $Y$ by observing $A$ while accounting for the entropy of the underlying distributions and adjusting chance based on the number of categories (as in [50, 51]). We rely on the faithfulness assumption which implies that if there is a relationship between $A,Y$, then $MI(A;Y)>0$.
For detectability, the objective is to determine the ability to infer the value $A$ from the data and we accomplish this by first training a surrogate model $f:A\rightarrow X$ to predict $A$ from $X$. We next estimate the utility and detectability between data and labels for both possible directions of dependency between data and labels. In many cases, clinician and domain expertise may allow us to eliminate certain directions of dependency thus improving the interpretability of the results and reducing overall computation. Our ultimate goal is to identify whether there is a relationship $A\leftrightarrow X$ and if so measure its strength. We consider the direction of dependence between data and labels in order to control for the potential of information leaking through $Y$ and this process relies on the assumed/known relationship between $Y,X$.
@@
In the causal case (i.e., $X\rightarrow Y$), we are able to directly estimate the relationship without conditioning. In fact, we cannot condition on $Y$ since $Y$ represents a collider (i.e., $Y$ may be dependent on both $X$ and $A$). In that case, measuring the strength of relationship between $X\leftrightarrow A$ given $Y$ could give falsely inflated results since conditioning on a collider $Y$ creates an otherwise non-existent association between its parents ($X,A$). For example, if two conditions both increase mortality rate through different pathways, then given that a patient has died, knowledge of either attribute provides information about the other (e.g., given $A$ is the presence or absence of trauma wounds and $X$ contains information relating to disease, a patient who died but did not have trauma wounds is more likely to have had disease: $A\leftrightarrow X|Y$ even though $A\perp X$ ). As a result, we do not condition for cases where we expect $X\rightarrow Y$ and instead directly estimate $MI(A;\hat{A})$.
@@ 4.4 Bounding Performance Risk
While utility and detectability measures provide a quantitative measure of bias, how to interpret them in the context of potential model performance risks is less clear. The magnitudes of these measures are not easily comparable across datasets so we cannot rely on thresholds or conventions to directly assess risk. We could instead look to the relative rankings of attributes with respect to utility/detectability to understand attribute risk ordering, but cannot directly translate values to drops in task-relevant metrics like AUC.
@@ 4.5.2 Stigmatizing Language in Electronic Health Records
We fine-tune BERT [10] models for both attribute prediction and each of the three clinical tasks: (1) compliance, (2) appearance and demeanor, and (3) credibility and obstinacy. Attributes we evaluate as potential shortcuts include patient race, sex, and clinical specialty (see Appendix A). The dataset consists of manually annotated samples with a context window of 10 words before and 10 after each identified potentially stigmatizing anchor word [21]. The full list of anchor words for each task is available in Appendix A.0.1. As each task has a unique set of anchors, the input and attribute metadata are different across tasks. Hyperparameters for all task and attribute models are held constant. We use AdamW [34] with a fixed learning rate of 5e-5 and weight decay of 1e-5, a batch size of 16, dropout with probability 0.1, 10 training epochs with early stopping, and class balanced cross-entropy loss for all experiments.
@@ 4.5.4 SPLIT method
As an alternate baseline form of estimating detectability and following [16], we test whether task model representations implicitly encode attribute information. Given a pre-trained classifier, we remove the final fully-connected layer and replace it with a randomly initialized linear layer. We freeze the weights of the pre-trained network and finetune only the linear layer to predict the specified attribute value. As before, we perform cross-validation and measure the performance of the finetuned model on the aggregated predictions across all folds. Better than chance performance is an indicator that model representations encode some degree of attribute information. However, it does not necessarily indicate that the dataset itself is biased as some attribute information may be relevant to solving the task even when the dataset is balanced with respect to the attribute itself.
@@ 5 Data Availability
Data for the ICU mortality prediction (https://mimic.mit.edu) and skin lesion classification cases (https://challenge.isic-archive.com/data) are publicly available. Because clinical notes from the Johns Hopkins Medicine (JHM) dataset contain identifiable information, they cannot be shared outside of our study team; however, all code to replicate the experiment on other datasets will be made available. Data was maintained, controlled and accessed only by Johns Hopkins University and not by the U.S. Food and Drug Administration.
6 Acknowledgment
MP acknowledges partial funding by appointment to the Research Participation Program at the Center for Devices and Radiological Health administered by the Oak Ridge Institute for Science and Education through an interagency agreement between the US Department of Energy and the US Food and Drug Administration. The mention of commercial products, their sources, or their use in connection with material reported herein is not to be construed as either an actual or implied endorsement of such products by the Department of Health and Human Services. This is a contribution of the U.S. Food and Drug Administration and is not subject to copyright